What method did you use to solve this, so I can solve the next one.
2y = x + 1
–2x – y = 7

Answer by mananth(16946)

-1 x + 2 y = 1 .............1
-2 x -1 y = 7 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 2
-1 x 2 y = 1
-4 x + -2 y = 14
Add the two equations
-5 x = 15
/ -5
x = -3
plug value of x in (1)
-1 x + 2 y = 1
3 + 2 y = 1
2 y = 1 -3
2 y = -2
y = -1

m.ananth@hotmail.ca